# Seeing Sound, Hearing Sight: Uncovering Modality Bias and Conflict of AI models in Sound Localization

### v1

@@ Abstract
Imagine hearing a dog bark and instinctively turning toward the sound—only to find a parked car, while a silent dog sits nearby. Such moments of sensory conflict challenge perception, yet humans flexibly resolve these discrepancies, prioritizing auditory cues over misleading visuals to accurately localize sounds. Despite the rapid advancement of multimodal AI models that integrate vision and sound, little is known about how these systems handle cross-modal conflicts or whether they favor one modality over another. Here, we systematicly and quantitatively investigate modality bias and conflict resolution in AI sound localization. We evaluate a wide range of state-of-the-art multimodal models and compare them against human performance in carefully designed psychophysics experiments spanning six audiovisual conditions, including congruent, conflicting, and absent visual and audio cues. Our results reveal that humans consistently outperform AI in sound localization and exhibit greater robustness to conflicting or absent visual information by effectively prioritizing auditory signals. In contrast, AI shows a pronounced bias toward vision, often failing to suppress irrelevant or conflicting visual input, leading to chance-level performance. To close this gap, we fine-tune a state-of-the-art multimodal model on our curated stereo audio-image dataset synthesized via 3D simulations. With fewer training examples, our resulting model outperforms existing AI trained on standard audio-visual benchmark datasets. Remarkably, consistent with neuroscience findings, our model exhibits a human-like horizontal localization bias—favoring left-right over up-down precision. This likely stems from training on stereo audio that mirrors the horizontal arrangement of human ears. These results highlight how the quality of sensory inputs and the physical structure of sensory systems jointly shape the precision of multimodal representation. All code, data, and models will be made public.
@@ 1 Introduction
Figure 1: (a) Sound source localization challenge in naturalistic images. The left shows a multi-source scenario with sounds from several objects (yellow boxes), while the right depicts a single-source case with only a telephone ring (red box). In both settings, the task is to localize all sounding objects based on two-channel stereo audio. To allow systematic and controllable benchmarking of human and AI performance, we focus on the single-source localization task (red), which remains challenging due to scene clutter, occlusions, and ambiguous visual cues. (b) Depth-aware stereo audio synthesis. In the 3D simulation environment, a human listener (interaural distance: 0.17m) is placed at the origin, facing the RGB image on the screen. This image and its depth image are aligned in the same direction. Using a spatial audio renderer and audio from our library, stereo sound of the target car (red box) in the RGB image can be synthesized (see Sec 2.1).
Sound source localization (SSL) is the task of identifying the spatial origin of a sound within a visual scene. It plays a fundamental role in both biological perception (van2019cortical, ) and artificial intelligence (AI) (jenni2023audio, ), enabling systems to connect what they hear with what they see. SSL requires integrating information across auditory and visual modalities to accurately determine not only what is sounding, but also where it is located. In neuroscience and psychology, SSL provides insights into how humans integrate complex multisensory information to direct attention (francl2022deep, ; oude2024triple, ) and make dynamic decisions (noppeney2021perceptual, ). In real-world practice, accurate SSL is essential for autonomous vehicles anticipating hazards (fang2022traffic, ), assistive devices guiding visually impaired users (dimas2021self, ), and interactive robots engaging in human-centered tasks (chen2025human, ; saeed2019multi, ; dworakowski2023robots, ). These applications require reliable detection of sound sources, particularly in cluttered or ambiguous scenes.
@@
Consider a real-life example in Fig. 1(a) depicting a busy traffic intersection: cars honk, people converse, birds chirp, and an ambulance siren approaches from behind a building. Humans must rapidly detect and localize the siren despite competing sounds, partial visual occlusion, and environmental noise. Such everyday scenes highlight the challenges of SSL: resolving ambiguous visual or auditory cues, handling conflicting signals across modalities, and focusing attention on the most relevant source amidst distractions. Building AI systems capable of performing SSL tasks in these conditions remains an open problem.
@@
Recent AI research has developed multimodal models for SSL (georgescu2023audiovisual, ; huang2023mavil, ; jia2025uni, ; li2021space, ; free, ). Methods such as contrastive learning (denseav, ; cavp, ; mo2022closer, ; qian2020multiple, ) aim to align audio-visual embeddings by maximizing the similarity between matching pairs and minimizing it between mismatched ones. Other approaches leverage cross-modal attention (mask, ; tar, ; lin2023unsupervised, ) in transformer architectures (is3, ; gao2024avsegformer, ; li2023catr, ) to allow sound features to dynamically query relevant visual regions. Some works (pv4, ; xuan2022proposal, ; senocak2022less, ) also incorporate object priors, leveraging knowledge of what typically makes sound to guide localization. However, despite these advances (gong2022contrastive, ), little is known about how such models behave when the modalities conflict or when biases emerge—such as favoring visual cues over auditory ones.
@@
To address this gap, we systematically evaluate model behavior under six controlled audio-visual conditions: (1) Congruent — audio and visual cues align in both semantics and location; (2) conflicting visual cues, where the visual scene misleads localization; (3) absent visual cues, where the sounding object is completely occluded in the visual scene; and (4) vision-only and (5) audio-only conditions, where either vision or audio is entirely omitted. Moreover, similar to the cocktail party problem (haykin2005cocktail, ; sun2024unveiling, ), we further extend the multi-instance SSL (wang2023robust, ) to single-source stereo audio, where multiple instances of the same semantic category are present, potentially distracting the model from correctly localizing the target object. These manipulations allow us to probe how models resolve cross-modal ambiguity and characterize their reliance on each modality. To benchmark these model behaviors, we additionally conduct human behavioral studies under the same experimental conditions. Results show a clear performance gap: humans consistently outperform AI models, particularly in handling both congruent and incongruent conditions.
@@
While numerous SSL datasets (lee2021acav100m, ; languagebind, ; ade20k, ) exist, they often suffer from limitations that hinder robust multimodal alignment in AI models. Often, these datasets (zhou2024audio, ; refavs, ; peavs, ; acappella, ; slakh, ; urbansed, ) consist of scenes with a single, large, centrally placed sounding object, making them vulnerable to visual shortcut learning, where models perform well without truly integrating audio information (cavp, ; is3, ; fnac, ). Others (vggss, ; is3, ) are constructed by pairing unrelated images and sounds from independent datasets (coco, ; imagenet, ; vggsound, ; FSDnoisy18k, ; DCASE, ; l3das22, ; vocalsound, ), leading to weak cross-modal entanglement. More recent efforts (sanguineti2021audio, ; fsd, ; aytar2016soundnet, ) synthesize audio based on physics-informed rules (mcdermott2011sound, ; gun, ) or generative AI models (zhou2020sep, ; lll, ), but they still frequently rely on mono audio, neglecting the richer spatial cues provided by stereo signals. Stereo audio, widely studied in neuroscience for its role in spatial hearing through inter-channel differences (woods2017headphone, ), remains underexplored in SSL benchmarks in AI.
@@
To address this gap, we introduce a method that uses 3D simulation engines to generate stereo audio for images by combining standalone image and sound datasets. Unlike neuroscience approaches that require physically setting up microphones in real-world spaces, which is both time-consuming and costly, our simulation-based method offers a scalable and efficient alternative for producing spatialized audio paired with complex visual scenes. With this method, we contribute a large-scale AudioCOCO dataset, comprising 1,003,747 image-audio pairs with ground truth annotations. By simulating stereo audio that adheres to physical principles of sound propagation and spatial cues, AudioCOCO provides realistic and diverse audio-visual scenes.
To further validate the utility of our dataset, we fine-tune a state-of-the-art (SOTA) AI model on stereo audio from our AudioCOCO dataset. The resulting model outperforms existing SOTA methods, which tend to be biased toward large, centered, or vocal objects (is3, ; cavp, ; refavs, ; avs, ). Neuroscience research francl2022deep shows that humans exhibit greater precision in horizontal (left-right) sound localization than vertical (up-down). Remarkably, despite zero training on any human data, the resulting model demonstrates this human-like trend which is not observed in existing AI systems. The stereo audio used to train our model is generated using 3D simulation engines and captures the interaural distance characteristic of human ears. These findings suggest that strong audio-visual alignment between AI and human behavior stems from both the quality of sensory inputs and the physical structure of human sensory systems. We highlight our key contributions below:
@@
1. We introduce a unified framework to systematically benchmark audio-visual localization models under modality conflicts, absence, and misalignment. We develop a scalable method to generate two-channel stereo audio for static images using 3D simulation engines, enabling the creation of diverse, naturalistic, and spatially grounded audio-visual datasets.
@@
2. We design and conduct psychophysics experiments to assess human strategies in resolving audio-visual conflicts, providing a strong baseline for AI-human comparison. We provide a detailed analysis of modality conflicts and biases in existing state-of-the-art SSL models and humans, highlighting key performance and behavioral differences under challenging multi-modal conditions.
3. We argue that strong audio-visual alignment between AI and human behavior emerges from high-quality sensory inputs and the anatomical structure of human sensory systems. To demonstrate this, we fine-tune a state-of-the-art model on our realistic, spatially grounded, well-controlled audio-visual dataset. The resulting model not only significantly outperforms existing AI systems but also displays human-like spatial biases in localization precision.
@@ 2.1 AudioCOCO dataset
Image selection. We adopt the MSCOCO image dataset coco for its wide coverage of everyday objects. To avoid overlap with data potentially seen during AI model training, we only use the 29,737 images from the test set. Since some objects are more likely to produce sound in real-world contexts, we manually selected 12 audible object categories from the MSCOCO annotations—spanning humans, animals, vehicles, and electronic devices. Recognizing that larger objects may be easier to detect and localize, we further categorize sounding objects by their relative size in the image. Object size is defined as the ratio of the object’s segmentation mask area to the total image area, independent of their real-world scale. We define three size bins: Size1 (0–5%), Size2 (5–15%), and Size3 (15–30%). Objects occupying more than 30% of the image are excluded, as they make the localization task trivially easy. To ensure class balance, we cap the number of images per object size per category at 150. For experimental control, each image contains only one target sounding object, with no other instances of the same category. For the multi-instance SSL task (see Sec 2.2), we select 150 images per object size per category, each containing 2–5 instances of the same semantic category, with exactly one designated as the sounding target. See Supp Fig. S1 and Supp Fig. S2 for the distribution of image counts by category and the spatial locations of target sounding objects.
@@
Audio selection. Prior work often pairs COCO images with external audio datasets like VGGSound vggsound or FSDnoisy18k FSDnoisy18k , but these audio sources often lack spatial and temporal consistency. For example, moving sound sources (e.g., a dog running across the video frames) cause spatial shifts that make them unsuitable for generating realistic stereo audio on static images. Additionally, recordings may include background noise or mixed sounds from multiple objects, reducing semantic clarity and overall quality. To address these issues, we apply the three filtering steps to the videos from VGGSound vggsound , ensuring that only high-quality, semantically aligned audios are retained. See Supp Sec. S1 and Supp Fig. S3 for detailed filtering steps and results. After applying these three filtering criteria, we have 720 high-quality audio clips from VGGSound vggsound . Combined with the 7,284 selected images from MSCOCO coco , we proceed to synthesize depth-aware stereo audio-image pairs.
@@
Depth-aware stereo audio synthesis. To generate spatialized stereo audio for their corresponding sounding objects on static images, we employ Unity bomatter2021pigs as our 3D simulation engine, which allows us to control precise object locations and simulate realistic stereo sound based on the spatial layout of visual scenes. As illustrated in Fig. 1(b), we define a Cartesian coordinate system within Unity. The listener is positioned at the origin (0, 0, 0). The computer screen displays the image to listener, is placed parallel to the x–z plane and aligned along the positive y-axis, with a fixed physical distance of 0.76 meters from the listener in the human psychophysics experiment (Sec 2.3). To estimate the depth of objects within the 2D image, we utilize the DepthAnything model depthanything , which outputs a relative depth map with values ranging from 0 to 10. However, without access to the original camera intrinsics of MSCOCO images, determining absolute scene scale in Unity is nontrivial. To address this, we normalize the depth values $d_{p}$ of the image as $d_{p}^{\text{norm}}=d_{p,\max}-d_{p}$ where $d_{p,\max}$ are the maximum depth values in the image. Sound loudness decreases logarithmically with distance from the listener. To maintain sound amplitudes within a comfortable and perceptible range, we rescale the normalized depth by 0.5. The final y-coordinate in Unity $y_{u}$ for the sounding object is computed as $y_{u}=d_{p}^{\text{norm}}/2+0.76$. For the x and z coordinates of the sounding object in Unity, we map the object’s pixel location from the 2D image to the physical dimensions of the monitor. Given that the display has a resolution of 90 pixels per inch, we compute $x_{u}=x_{p}/90$ and $z_{u}=z_{p}/90$, where $x_{p}$ and $z_{p}$ are the pixel coordinates of the target object’s center in the image. Once the sounding object’s 3D position $(x_{u},y_{u},z_{u})$ is established in Unity, we place a static audio source at this location. Using an interaural distance of 0.17 meters, reflecting the typical distance between human ears, Unity simulates realistic two-channel stereo sound based on the spatial relationship between the listener and the sound source. This procedure allows us to synthesize spatially grounded, depth-aware stereo audio for each image-sound pair.
2.2 Experimental conditions
As illustrated in Fig. 2, we introduce six experimental conditions to systematically examine the modalitity biases and conflicts in humans and AI models. Each condition comprises 184,227 image-audio pairs, except for the VOnly and MultiInstLoc conditions.

### v2

@@ Abstract
Imagine hearing a dog bark and instinctively turning toward the sound—only to find a parked car, while a silent dog sits nearby. Such moments of sensory conflict challenge perception, yet humans flexibly resolve these discrepancies, prioritizing auditory cues over misleading visuals to accurately localize sounds. Despite the rapid advancement of multimodal AI models that integrate vision and sound, little is known about how these systems handle cross-modal conflicts or whether they favor one modality over another. Here, we systematically and quantitatively examine modality bias and conflict resolution in AI models for Sound Source Localization (SSL). We evaluate a wide range of state-of-the-art multimodal models and compare them against human performance in psychophysics experiments spanning six audiovisual conditions, including congruent, conflicting, and absent visual and audio cues. Our results reveal that humans consistently outperform AI in SSL and exhibit greater robustness to conflicting or absent visual information by effectively prioritizing auditory signals. In contrast, AI shows a pronounced bias toward vision, often failing to suppress irrelevant or conflicting visual input, leading to chance-level performance. To bridge this gap, we present EchoPin, a neuroscience-inspired multimodal model for SSL that emulates human auditory perception. The model is trained on our carefully curated AudioCOCO dataset, in which stereo audio signals are first rendered using a physics-based 3D simulator, then filtered with Head-Related Transfer Functions (HRTFs) to capture pinnae, head, and torso effects, and finally transformed into cochleagram representations that mimic cochlear processing. To eliminate existing biases in standard benchmark datasets, we carefully controlled the vocal object sizes, semantics, and spatial locations in the corresponding images of AudioCOCO. EchoPin outperforms existing models trained on standard audio-visual datasets. Remarkably, consistent with neuroscience findings, it exhibits a human-like localization bias, favoring horizontal (left–right) precision over vertical (up–down) precision. This asymmetry likely arises from HRTF-shaped and cochlear-modulated stereo audio and the lateral placement of human ears, highlighting how sensory input quality and physical structure jointly shape precision of multimodal representations. All code, data, and models are available here.
@@ 1 Introduction
Figure 1: (a) Sound source localization challenge in naturalistic images. In the multi-source scenario (orange panel), multiple objects in the scene—such as pedestrians, birds, cars, and trucks (highlighted by yellow boxes)—emit sounds, whereas in the single-source scenario (red panel), only one object, a telephone (red box), produces sound. In both settings, the task is to localize all sounding objects based on two-channel stereo audio. To allow systematic and controllable benchmarking of human and AI performance, we focus on the single-source localization task (red panel), which remains challenging due to scene clutter, occlusions, and ambiguous visual cues. (b) Depth-aware stereo audio synthesis. In the 3D simulator, a human listener (interaural distance: 0.17m) is placed at the origin, facing the RGB image on the screen. This image and its depth image are aligned in the same direction. Using a spatial audio renderer and audio from our library, stereo audio of the target car (red box) in the RGB image can be synthesized (see Sec 2.1).
@@
Sound source localization (SSL) is the task of identifying the spatial origin of a sound within a visual scene. It plays a fundamental role in both biological perception (van2019cortical, ) and artificial intelligence (AI) (jenni2023audio, ), enabling systems to connect what they hear with what they see. In natural environments, visual inputs interact with auditory cues, often dominating or recalibrating sound perception—a phenomenon exemplified by classic cross-modal illusions (noppeney2021perceptual, ). In practice, accurate SSL is critical for real-world systems such as autonomous vehicles anticipating hazards (fang2022traffic, ), assistive technologies for visually impaired users (dimas2021self, ), and robots operating in human-centered environments (chen2025human, ; saeed2019multi, ; dworakowski2023robots, ; li2022learning, ; zhu2019deformable, ). These applications demand robust auditory inference under noisy, cluttered, and ambiguous sensory conditions. Consider a real-life example in Fig. 1(a) depicting a busy traffic intersection: cars honk, people converse, birds chirp, and an ambulance siren approaches from behind a building. Humans must rapidly detect and localize the siren despite competing sounds, partial visual occlusion, and environmental noise. Such everyday scenes highlight the challenges of SSL: resolving ambiguous visual or auditory cues, handling conflicting signals across modalities, and focusing attention on the most relevant source amidst distractions. Building AI systems capable of performing SSL tasks in these conditions remains an open problem.
@@
Recent AI research has developed multimodal models for SSL (georgescu2023audiovisual, ; huang2023mavil, ; jia2025uni, ; li2021space, ; free, ; jenni2023audio, ; senocak2025toward, ). Methods such as contrastive learning (denseav, ; mo2022closer, ; qian2020multiple, ) aim to align audio-visual embeddings by maximizing the similarity between matching pairs and minimizing it between mismatched ones. Other approaches leverage cross-modal attention (mask, ; tar, ; lin2023unsupervised, ) in transformer architectures (gao2024avsegformer, ; li2023catr, ) to allow sound features to dynamically query relevant visual regions. Some works (pv4, ; xuan2022proposal, ; senocak2022less, ) also incorporate object priors, leveraging knowledge of what typically makes sound to guide localization. However, despite these advances (gong2022contrastive, ), little is known about how such models behave when the modalities conflict or when biases emerge—such as favoring visual cues over auditory ones.
@@
To address this gap, we systematically evaluate model behavior under six controlled audio-visual conditions: (1) Congruent — audio and visual cues align in both semantics and location; (2) conflicting visual cues, where the visual scene misleads localization; (3) absent visual cues, where the sounding object is completely occluded in the visual scene; and (4) vision-only and (5) audio-only conditions, where either vision or audio is entirely omitted. Moreover, similar to the cocktail party problem (haykin2005cocktail, ; sun2024unveiling, ), we further extend the single-source stereo audio to (6) the multi-instance SSL (wang2023robust, ), where multiple instances of the same semantic category are present, potentially distracting the model from correctly localizing the target instance. These manipulations allow us to probe how models resolve cross-modal ambiguity and characterize their reliance on each modality. To benchmark these model behaviors, we additionally conduct human behavioral studies under the same experimental conditions. Results show a clear performance gap: humans consistently outperform AI models in handling both congruent and incongruent conditions.
@@
While numerous SSL datasets (lee2021acav100m, ; languagebind, ; ade20k, ) exist, they often suffer from limitations that hinder robust multimodal alignment in AI models. Typically, these datasets (zhou2024audio, ; refavs, ; avs, ; peavs, ; acappella, ; slakh, ; urbansed, ) consist of scenes with a single, large, centrally placed sounding object, making them vulnerable to visual shortcut learning, where models perform well without truly integrating audio information (cavp, ; fnac, ). Others (vggss, ; is3, ; torralba2011unbiased, ) are constructed by pairing unrelated images and sounds from independent datasets (coco, ; imagenet, ; vggsound, ; FSDnoisy18k, ; DCASE, ; l3das22, ; vocalsound, ; politis2022starss22, ; lollmann2018locata, ), leading to weak cross-modal entanglement. More recent efforts (sanguineti2021audio, ; fsd, ; aytar2016soundnet, ) synthesize audio based on physics-informed rules (mcdermott2011sound, ; gun, ) or generative AI models (zhou2020sep, ; lll, ), but they still frequently rely on mono audio, neglecting the richer spatial cues provided by stereo audio signals.
@@
Inspired by neuroscience findings highlighting the role of inter-channel differences in spatial hearing (woods2017headphone, ), we propose a method that leverages 3D simulation engines to generate stereo audio from images by integrating separate image and sound datasets. Unlike neuroscience approaches that require physically setting up microphones in real-world spaces, which is both time-consuming and costly, our simulation-based method offers a scalable and efficient alternative for producing spatialized audio paired with complex visual scenes. With this method, we contribute a large-scale AudioCOCO dataset, comprising 28,224 image-audio pairs with ground truth annotations. By simulating stereo audio that adheres to physical principles of sound propagation and spatial cues, AudioCOCO provides realistic and diverse audio-visual scenes.
Human sound localization relies on a cascade of acoustic transformations shaped by the ear, head, and torso. Direction-dependent spectral notches introduced by the pinna help resolve elevation and front–back ambiguities (van2019cortical, ; zilany2009phenomenological, ; mcwalter2019illusory, ). These cues are formalized in the Head-Related Transfer Function (HRTF), which encodes interaural time (ITD) and level differences (ILD) alongside fine-grained spectral features. At the cochlea, sounds are further decomposed into frequency-selective channels that preserve ITD/ILD while transforming HRTF-induced modulations into neural representations.
To model these biological mechanisms, we introduce EchoPin, a neuroscience-inspired model for SSL. EchoPin pre-processes stereo audio using Head-Related Transfer Function (HRTF)–based filtering and cochleagram representations derived from Equivalent Rectangular Bandwidth (ERB) filters. These designs capture the tonotopic organization and temporal dynamics of the auditory periphery (saddler2021deep, ; saddler2024models, ) more faithfully than conventional mel-spectrograms (shimada2025stereo, ; yu2025two, ; mu2024seld, ). EchoPin then employs dual encoders to jointly process audio and visual inputs, trained with contrastive learning on our AudioCOCO dataset. Experimental results suggest that EchoPin outperforms existing models. Notably, without any human behavioral supervision, EchoPin reproduces a human-like horizontal localization bias (francl2022deep, ), an emergent property that was not prominent in previous AI systems. We attribute this to EchoPin’s 3D stereo audio pipeline, which integrates interaural spacing, HRTF filtering, and ERB-based cochlear processing to jointly enhance sensory fidelity and multimodal alignment. We highlight our key contributions below:
@@
1. We introduce a unified framework to systematically benchmark audio-visual localization models under modality conflicts, absence, and misalignment. We propose a scalable pipeline that synthesizes two-channel stereo audio for static images via 3D simulation, and construct a large-scale, naturalistic, and spatially grounded audio-visual dataset, named as AudioCOCO.
@@
2. We design and conduct psychophysics experiments to assess human strategies in resolving audio-visual conflicts, providing a strong baseline for AI-human comparison. We provide a detailed analysis of modality conflicts and biases in existing SSL models and humans, highlighting key performance and behavioral differences under challenging multimodal conditions.
3. We introduce EchoPin, a neuroscience-inspired model trained on our curated AudioCOCO dataset. It features a dual-encoder architecture that processes stereo audio–visual pairs. The stereo audio is pre-processed using HRTF-based spatial filtering and cochlear-inspired frequency decomposition. Experimental results show that precise audio–visual alignment emerges from high-fidelity sensory inputs and biologically grounded ear-structure priors. EchoPin not only achieves superior localization accuracy but also exhibits human-like localization biases, favoring horizontal over vertical precision.
@@ 2.1 AudioCOCO dataset
Image selection. We use the MSCOCO coco dataset for its broad coverage of everyday objects and apply our selection criteria below to all images from the standard training and test sets. From its dataset annotations, we manually select 12 audible object categories encompassing humans, animals, vehicles, and electronic devices. Recognizing that larger objects may be easier to detect and localize, we further categorize sounding objects by their relative size in the image. Object size is defined as the ratio of the object’s segmentation mask area to the total image area, independent of their real-world scale. We define three size bins: Size1 (0–5%), Size2 (5–15%), and Size3 (15–30%). Objects occupying more than 30% of the image are excluded, as they make the localization task trivially easy. To maintain class balance, we limit the number of images per object size per category to 150 within each training or test set. For experimental control, each image contains only one target sounding object, without any other instances of the same category. For the multi-instance SSL (Sec 2.2), we select 150 images per object size per category, each containing 2–5 instances of the same semantic category, with exactly one designated as the sounding target. See Supp Fig. S1, Supp Fig. S2, and Supp Fig. S3 for the distribution of image counts by category and their target spatial locations in both the training and test sets. After applying the selection criteria, we randomly sample 4,953 qualified images from the MSCOCO training set for training, and 5,500 images from the official test set—comprising 2,840 single-object and 2,660 multi-object scenes—for testing.
@@
Audio selection. Prior work often pairs MSCOCO images with external audio datasets like VGGSound vggsound or FSDnoisy18k FSDnoisy18k , but these audio sources often lack spatial and temporal consistency. For example, moving sound sources (e.g., a dog running across the video frames) cause spatial shifts that make them unsuitable for generating realistic stereo audio on static images. Additionally, recordings may include background noise or mixed sounds from multiple objects, reducing semantic clarity and overall quality. To address these issues, we apply three filtering steps to the VGGSound videos, retaining 3,533 clips from the standard training set and 727 clips from the standard test set that contain high-quality audio, spatially and semantically aligned with the visual content. See Supp Sec. S1, Supp Fig. S4, and Supp Fig. S5 for detailed filtering steps and results.
Next, we randomly pair the selected MSCOCO images with high-quality audios from VGGSound to construct the AudioCOCO dataset. The dataset comprises 9,360 audio–image pairs in the training set and 18,864 pairs in the test set, which are further divided into six experimental conditions (Sec. 2.2).
@@
Depth-aware stereo audio synthesis. To generate spatialized stereo audio for their corresponding sounding objects on static images, we employ Unity bomatter2021pigs as our 3D simulation engine, which allows us to control precise object locations and simulate realistic stereo sound based on the spatial layout of visual scenes. As illustrated in Fig. 1(b), we define a Cartesian coordinate system within Unity. The listener is positioned at the origin (0, 0, 0). The computer screen displays the image to listener, is placed parallel to the x–z plane and aligned along the positive y-axis, with a fixed physical distance of 0.76 meters from the listener in the human psychophysics experiment (Sec 2.3). To estimate the depth of objects within the 2D image, we utilize the DepthAnything model depthanything , which outputs a relative depth map with values ranging from 0 to 10. However, without access to the original camera intrinsics of MSCOCO images, determining absolute scene scale in Unity is nontrivial. To address this, we normalize the depth values $d_{p}$ of the image as $d_{p}^{\text{norm}}=d_{p,\max}-d_{p}$ where $d_{p,\max}$ are the maximum depth values in the image. Sound loudness decreases logarithmically with distance from the listener. To maintain sound amplitudes within a comfortable and perceptible range, we rescale the normalized depth by 0.5. The final y-coordinate in Unity $y_{u}$ for the sounding object is computed as $y_{u}=d_{p}^{\text{norm}}/2+0.76$. For the x and z coordinates of the sounding object in Unity, we map the object’s pixel location from the 2D image to the physical dimensions of the monitor. Given that the display has a resolution of 90 pixels per inch, we compute $x_{u}=x_{p}/90$ and $z_{u}=z_{p}/90$, where $x_{p}$ and $z_{p}$ are the pixel coordinates of the target object’s center in the image.
Once the sounding object’s 3D position $(x_{u},y_{u},z_{u})$ is established in Unity, we place a static audio source at this location. Using an interaural distance of 0.17 meters, reflecting the typical distance between human ears, Unity simulates realistic two-channel stereo sound based on the spatial relationship between the listener and the sound source. This procedure allows us to synthesize spatially grounded, depth-aware stereo audio for each image-sound pair.
2.2 Experimental conditions in the AudioCOCO test set
As shown in Fig. 2, the AudioCOCO test set includes six experimental conditions, totaling 18,864 image–audio pairs to systematically probe modality biases and conflicts in humans and AI models. Each condition contains 2,900 pairs, except MultiInstLoc, which includes 4,364. All models are trained only on congruent conditions to evaluate generalization, while these six conditions are used exclusively for testing.